OpenStudy (xapproachesinfinity):

hey i had a disagreement today with my prof we were doing l'hopital's rule with \[\lim_{x\to \infty}(1+x)^{\frac{1}{x}}\] I said that the limit should be \[x \rightarrow o^{+}\] not infinity in order for us to get e what he did lead to \[e^{\lim_{x\to \infty}\frac{1}{1+x}}\] which i said goes to 1 not e

OpenStudy (rational):

yes whats the exact question ?

OpenStudy (xapproachesinfinity):

as i said my prof said that that limit gives e which is not true for me! i disagreed with him

OpenStudy (anonymous):

You're right, I believe your professor meant one of the following: \[\lim_{x\to\infty}\left(1+\frac{1}{x}\right)^x=\lim_{x\to0}(1+x)^{1/x}=e\]

OpenStudy (xapproachesinfinity):

yeah exactly!